The cost function of a monopoly firm is TC = 10 + 2Q where TCq is the total cost of producing the quantity. The demand in this market is expressed by the equation Q 14 = 14 – P where P denotes the price. Estimate the profit to be made by the monopolist.

1
Expert's answer
2022-02-08T11:43:53-0500

"Profit= TR-TC"

"TC=10+2Q"

From the demand curve:

"Q=1-0.71P" f

"TR=(1-0.071P)\times{P}=P-0.071P^{2}"

"Profits=(P-0.071P^{2})-(10-2Q)"

"=(p-0.071P^{2})-(10+2-0.142P)"

"=-0.071P^{2}+1.143P-12"

Using the quadratic formula to solve for P:

"P=\frac{-b+-√(b^{2}-4ac)}{2a}"

"P=\frac{-1.143+-√(1.143^{2}-4×-0.071×12)}{2×-0.071}"

"P=6.91"

"Q=1-0.071\u00d76.91=0.51"

Therefore

"TR=6.91\u00d70.51=3.52"

"TC=10+2\u00d70.51=11.02"

"Profit=3.52-11.02=-7.5"

Therefore the firm is making a loss.